OpenStudy (anonymous):

Let f(x) = (1-t^2)^1/3 from the integral 1 to cosx. Find and simplify f'(x)

OpenStudy (anonymous):

would it be correct to replace t with cos x and then multiply with derivative?

OpenStudy (anonymous):

or am I mistaking that process for if it was integral 0 to cos x

OpenStudy (anonymous):

Use the fundamental theorem of calculus: \[\frac{d}{dx}\left[\int_c^{g(x)}f(t)~dt\right]=f(g(x))\cdot g'(x)\] where \(c\) is a constant.

OpenStudy (anonymous):

oh gotcha tysm